brooklynp:

what is the square root of 175

velmalovesshaggy145:

You'd get a number with a decimal I know that much

velmalovesshaggy145:

13^2 is 169 right

velmalovesshaggy145:

right

velmalovesshaggy145:

I can't give you a direct answer or I could get reported/or given a warning

velmalovesshaggy145:

and 14^2 is 196 So the decimal would be after 13

velmalovesshaggy145:

But basically rounded you get, 13.23

brooklynp:

thanks cool

velmalovesshaggy145:

Medal

velmalovesshaggy145:

medal please*

Luigi0210:

Well, if you want an exact answer, you can just break up the number: \[\large \sqrt{175~} \] is the same as \[\large \sqrt{(25)(7)~} \] And I think you know that the square root of 25 is 5, so you can pull out a five and are left with your exact answer of: \[\Large 5\sqrt{7~}\]

velmalovesshaggy145:

@4udr1 Luigi killed that answer, but can you give me a medal too